OmniBDE has three major competitors: RR Donnelly (P/E = 6.0), CDK Global (P/E = 27.7), and Interpublic (P/E = 14.0) . OmniBDE current EPS is expected to be $5.70. If OMC’s current price is 75.57, is it over or under valued?
Answer #1

Given EPS=5.70

current price=75.57

Therefore, P/E ratio= 75.57/5.70=13.25

Now if we look at the industry P/E Ratios of Competitors its 6, 27.7and 14 respectively.

Hence if we take the average its equal to 15.9 which is above the P/E ratio of OMC. Hence OMC is undervalued.